# FastVGGT: Training-Free Acceleration of Visual Geometry Transformer

### v1

@@ 5 Conclusion
In this work, we proposed Fast-VGGT, a training-free method that accelerates VGGT inference via strategic token merging without sacrificing reconstruction quality. By profiling VGGT, we identified the global attention module as the primary bottleneck for long-sequence inputs and observed strong similarities in its attention maps. To address this, we introduce token merging to alleviate the inference bottleneck and design strategies tailored to the feed-forward visual geometry model. Experiments on multiple benchmarks show that FastVGGT achieves up to 4x speedup on 1000-images inputs while maintaining competitive accuracy in both camera pose estimation and 3D reconstruction tasks, and further mitigates error accumulation in long-sequence settings. These results highlight token merging as a principled solution for scaling visual geometry models and demonstrate FastVGGT’s immediate practicality for real-world applications

### v2

@@ 5 Conclusion
In this work, we proposed FastVGGT, a training-free method that accelerates VGGT inference via strategic token merging without sacrificing reconstruction quality. By profiling VGGT, we identified the global attention module as the primary bottleneck for long-sequence inputs and observed strong similarities in its attention maps. To address this, we introduce token merging to alleviate the inference bottleneck and design strategies tailored to the feed-forward visual geometry model. Experiments on multiple benchmarks show that FastVGGT achieves up to 4x speedup on 1000-images inputs while maintaining competitive accuracy in both camera pose estimation and 3D reconstruction tasks, and further mitigates error accumulation in long-sequence settings. These results highlight token merging as a principled solution for scaling visual geometry models and demonstrate FastVGGT’s immediate practicality for real-world applications.